It takes a boat 4 hours to travel 24 miles down a river and 6 hours to return upstream to its starting point. What is the rate of the current in the river?
Answer by checkley77(12569)

24/4=6 mph downstream.
24/6=4 mph upstream.
(6-4)/2=2/2=1 is the speed of the current.